C=200+0.5(Y-T)
T=100+0.2Y
Nx=150-0.2Y
I=300
G=400

a) What is the value of government budget deficit/surplus?

b) Assume that the government introduces an insurance of 2% to the households show this will affect consumption and determine the new equilibrium level of income.
1
Expert's answer
2020-06-29T17:31:37-0400

a) Value of government budget deficit/surplus

National Income = Consumption Expenditure + Investment + Government Expenditure + Net exports

Y = C+I+G+NX

Y = 200+0.5 (Y-T) + 300 + 400 + 150-0.2Y

T=100+0.2Y

Y = 0.5 (Y- (100+0.2Y)) + 1050 - 0.2Y

Y = 0.5 (0.8Y- 100) + 1050 - 0.2Y

Y = 0.4Y- 50 + 1050 - 0.2Y

Y – 0.2Y = 1000

Y = 1000 /0.8

Y = 1250

Budget deficit/surplus = G – T

Government Expenditure = 400

Tax income = T=100+0.2Y

T=100+0.2(1250)

T = 100 + 250

T = 350

Budget deficit = 400 – 350

Budget deficit = 50

b) Effect of introduction of insurance of 2% to the households and the new equilibrium level of income.

Insurance = 2%

Consumption expenditure

C=200+0.5(Y-T)

C=200+0.5(1250 - 350)

C=200+0.5(900)

C=200+ 450

C= 650

Introduction of insurance of 2% increases the consumption expenditure

Insurance = "2\%\times650"

Insurance = 13

New equilibrium level of income = Y = C+I+G+NX

Y = (650+13) + 300 + 400 + 150 - 250

Y = 1,263